Find the coordinates of the foci of the hyperbola: x^2/16 - y^2/4 = 1
Answer by Nate(3500)

x^2/16 - y^2/4 = 1
Transverse Axis: horizontal
Center: (0,0)
a = 4
b = 2
c^2 = a^2 + b^2
c = sqrt(16 + 4)
c = sqrt(20) = 2*sqrt(5)
Vertices: (-4,0) and (4,0)
Foci: (-2*sqrt(5),0) and (2*sqrt(5),0)